OpenStudy (anonymous):

Now how am i suppose to do this Phil hit 4 more home runs than Charlie did last season. The product of their home runs from last season is 285. How many home runs did Charlie hit last season? 281 23 19 15

OpenStudy (anonymous):

its b or c

OpenStudy (anonymous):

it would be D

OpenStudy (anonymous):

nor b or c because if you Let x represent the number of home runs Charlie hit last season. This means that the number of home runs that Phil hit last season is represented by x + 4. Since the product of their home runs is 285, the following equation can be used to solve for x.

OpenStudy (jojo4eva):

Yeh the answer is D. All you have to do is do 285/15 and you get 19 so your answer are 15 and 19

OpenStudy (jojo4eva):

so the answer is D

DivineSolar (divinesolar):

Yes it is D.

OpenStudy (anonymous):

sooooo x(x + 4) =285 x^2 + 4x =285 x^2 + 4x - 285 =0 (x + 19)(x - 15) =0 x + 19=0 x - 15=0 x=-19 x=15

OpenStudy (anonymous):

Since it is impossible to hit a negative number of home runs, Charlie hit 15 home runs

OpenStudy (anonymous):

lol im a genius huh

OpenStudy (jojo4eva):

So conceited hun lol

OpenStudy (jojo4eva):

Im still weak off how you solved your own question and got a medal for it lol im laughing so hard right now

OpenStudy (anonymous):

lol what i was baffled at first but got a little help of an explanation and just killed it @jojo4eva

OpenStudy (jojo4eva):

lol ayye thats whatsup lol you better than me haha

OpenStudy (anonymous):

*popping my hoodie pajamas*

OpenStudy (jojo4eva):

haha im dead lol